The equilibrium constant, Kc, for the following reaction is 5.89×10-4 at 635 K.

NH4I(s)<-->NH3(g) + HI(g)


Calculate Kc at this temperature for the following reaction:

NH3(g) + HI(g) <-->NH4I(s)

Kc=?
Answer #1

The given reaction is:

NH4I(s) <-> NH3(g) + HI(g)

Kc = [NH3][HI]

= 5.89*10^-4

The required reaction is:

NH3(g) + HI(g) <-> NH4I(s)

Use:

Kc’ = 1/([NH3][HI])

= 1/ Kc

= 1/(5.89*10^-4)

= 1.70*10^3

Answer: 1.70*10^3